OpenStudy (anonymous):

What is the slope of the line between (–4, 4) and (–1, –2)?

OpenStudy (mertsj):

Subtract the y values. Subtract the x values. Put the first number over the second number. That is the slope. (Make sure you subtract the y's and x's in the same order.

OpenStudy (ash2326):

deesmoopth you know slope's formula?

OpenStudy (anonymous):

\[y _{2}-y _{1}=m(x _{2}-x _{1})\] where m is the slope.

OpenStudy (anonymous):

slope=4-(-2)/-4-(-1) =6/-3 =-2

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

sorry?